OpenStudy (anonymous):

find the coefficient of x^6 in the expansion of (2x+3)^10

OpenStudy (anonymous):

start by multiplying (2x+3) by itself 10 times. add up all the coefficients of like terms. see what the coefficient of x^6 is

OpenStudy (anonymous):

thanks! (:

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

what is (2x+3) by itself ten times

OpenStudy (anonymous):

(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)(2x+3)

OpenStudy (anonymous):

I think you can be a little less crude about it, and simply write down: \[ \binom{10}{6} \times 2^6 \times 3^4 \]

OpenStudy (anonymous):

Where \[\binom{n}{r} = ^nC_r = \frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

Newton is right. its much easier to do it his way.

OpenStudy (anonymous):

i have another question